OpenStudy (anonymous):

sqrt{5x+29=x+3} sqrt{5x+29=x+3} @Mathematics

OpenStudy (anonymous):

\[\sqrt{5x+29=x+3}\]

OpenStudy (anonymous):

x=-6.5

jhonyy9 (jhonyy9):

so than sqrt(5x+29 = sqrt(x+3) squared both sides 5x+29=x+3 4x=-26 2x= -13 x= -13/2